OpenStudy (anonymous):

show that (gr o gr^-1) (x) = x for all x ∈ D gr gr = x^2+5x-6 gr^-1- (x+12.25)^0.5 -2.5

OpenStudy (anonymous):

gr(x)=x^2+5X-6 ,gr^-1(x)=-5/2 +sqrt(x+1/4) gr(-5/2+sqrt(x+1/4)=(-5/2+sqrt(x+1/4))^2+5(-5/2+sqrt(x+1/4)-6

OpenStudy (anonymous):

\[(\sqrt{x+12.25}-2.5)^2+5(\sqrt{x+12.25}-2.5)-6\rightarrow \]

OpenStudy (anonymous):

\[(\sqrt{x+12.25})^2-2*2.5(\sqrt{x+12.25})+2.5^2+5(\sqrt{x+12.25})-2.5*5-6\]

OpenStudy (anonymous):

knock were did you get ( squrt x+12.25 -2.5 ? the squart?

OpenStudy (anonymous):

\[x+12.25+2.5^2-2.5*5-6=x\]

OpenStudy (anonymous):

knock were did you get the squart

OpenStudy (anonymous):

got it thanks

OpenStudy (anonymous):

\[\sqrt{x+12.25}=(x+12.25)^\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

(: